OpenStudy (anonymous):

For a given arithmetic sequence, the 42nd (a)42=286 and 91st term =629 find the value of the 12th term (a)12

OpenStudy (fibonaccichick666):

well, that's a crumby question. ew. ok, so let's think for a moment, what have you tried?

Nnesha (nnesha):

ikr! i also want to know how to solve that ;

OpenStudy (fibonaccichick666):

or, what properties do you know about an arithmatic sequence?

OpenStudy (anonymous):

You can first find the difference between each term by subtracting 286 from 629. That'll give you the space between 91st and 42nd term. You then divide the answer, 343, by 49, which are then terms between the 91st and 42nd. Your pattern therefore would be +7 increase per term. So, find term one by working backwards. You'd have to go back 41 terms to get back to 1, so 41 times 7 equals 287. 286-287=-1 To find the 12th term, you'd have to use the equation for arithmetic formulas, which is An=A1+(n-1)d. A12=A1+(12-1)7 Simplify, and you get A12=A1+77. So, A12=-1+77 So, A12= 76